OpenStudy (anonymous):

The solution to 2x - 2 + 5 = 13 is _____.

OpenStudy (anonymous):

2x+3=13 2x=10 x=10/2 x=5

OpenStudy (anonymous):

2x+3=13 2x=10 x=10/2 x=5

OpenStudy (anonymous):

The sum of five consecutive integers is 75. What is the third number? 13 14 15 16

OpenStudy (anonymous):

Solve x + x + 2 + 2x = -2. -2 -1 0 1